OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 7. The perimeter of the larger rectangle is 42 inches. What is the perimeter of the smaller rectangle? A. 10.5 inches B. 73.5 inches C. 28 inches D. 24 inches

OpenStudy (anonymous):

It's c?

OpenStudy (xapproachesinfinity):

let's name the perimeter of the small x and the large y so \[\frac{y}{x}=4/7\] they gave us one perimeter y=42 so \[\frac{42}{x}=4/7\]

OpenStudy (anonymous):

42/7 = 6 x4 = 24

OpenStudy (calculusxy):

This is a proportion of smaller to larger: 4. x -- = --- 7. 42

OpenStudy (anonymous):

Oh...D...Sorry. I read it wrong. Thank you. All of you.

OpenStudy (calculusxy):

So just solve for x.

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (xapproachesinfinity):

i turned out i wrote it the wrong way should be smaller to larger

OpenStudy (anonymous):

It's okey

OpenStudy (xapproachesinfinity):

but you got the point :)

OpenStudy (anonymous):

I passed C:

OpenStudy (anonymous):

I say D